OpenStudy (anonymous):

Best answer? Will give medal. I've been looking for an explanation to this question for a while. Anybody know how to get the answer? Solve the compound inequality. Write solution as a compound inequality. -4/3 is less than or equal to 1/7 w-3/4<1

OpenStudy (solomonzelman):

1/7<w-3/4<1 this?

OpenStudy (solomonzelman):

can you draw it?, don't exactly get the question!

OpenStudy (anonymous):

-4/3 <= 1/7 w - 3/4 <1

OpenStudy (solomonzelman):

\[-4/3 <= 1/7 w - 3/4 <1\]

OpenStudy (solomonzelman):

add 3/4 to every part of inequility, to -4/3 to 1/7 w - 3/4 & to 1

OpenStudy (anonymous):

-7/12 <= 1/7 w < 1 3/4

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

-4/3 <= 1/7 w - 3/4 <1 \[-4/3 + 3/4<= 1/7 w - 3/4 + 3/4<1+ 3/4\] what do you get now?

OpenStudy (solomonzelman):

I think you are right! I didn't actually count this!

OpenStudy (solomonzelman):

Yes you are right, !

OpenStudy (anonymous):

Yay!

OpenStudy (solomonzelman):

next step, everything times 7.

OpenStudy (anonymous):

-4 1/12 <= 1w < 12 1/4

OpenStudy (anonymous):

Not sure about the middle part though...

OpenStudy (solomonzelman):

This is right!

OpenStudy (anonymous):

Good. :D

OpenStudy (anonymous):

What's next?

OpenStudy (solomonzelman):

Nothing is next 1w=w right? So this is your answer \[-4 1/12 <= w < 12 1/4\] unlesss you need to graph it, do you?

OpenStudy (solomonzelman):

"Write solution as a compound inequality." and you just did!

OpenStudy (anonymous):

Oh...

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

@DelilahRobo, need more help?

OpenStudy (anonymous):

Uhm. Write the solution of the inequality in set-builder notation. 2(3p - 11) >= -16

OpenStudy (solomonzelman):

I've never done anything in a set builder notation. Weird, but sorry...! You can make a new question, and tag someone else.

OpenStudy (anonymous):

That's okay. Thanks a lot for the help. I have been trying to figure out that first question forever!